# MRGen: Diffusion-based Controllable Data Engine for MRI Segmentation towards Unannotated Modalities

### v1

@@ Abstract
Medical image segmentation has recently demonstrated impressive progress with deep neural networks, yet the heterogeneous modalities and scarcity of mask annotations limit the development of segmentation models on unannotated modalities. This paper investigates a new paradigm for leveraging generative models in medical applications: controllably synthesizing data for unannotated modalities, without requiring registered data pairs. Specifically, we make the following contributions in this paper: (i) we collect and curate a large-scale radiology image-text dataset, MedGen-1M, comprising modality labels, attributes, region, and organ information, along with a subset of organ mask annotations, to support research in controllable medical image generation; (ii) we propose a diffusion-based data engine, termed MRGen, which enables generation conditioned on text prompts and masks, synthesizing MR images for diverse modalities lacking mask annotations, to train segmentation models on unannotated modalities; (iii) we conduct extensive experiments across various modalities, illustrating that our data engine can effectively synthesize training samples and extend MRI segmentation towards unannotated modalities. The code, model, and data will be publicly available at https://haoningwu3639.github.io/MRGen/.
Figure 1: Motivations and Overview. Left: The heterogeneity of MRI modalities poses challenges to the generalization of segmentation models. Our proposed data engine, MRGen, addresses this by controllably synthesizing training data, enabling segmentation towards mask-unannotated modalities. Right: (a) Previous generative models rely on mask-annotated data and are restricted to data augmention in training modalities; (b) Image translation typically requires registered data pairs (indicated by dashed lines), limited to conversions between specific modalities; (c) MRGen introduces a new paradigm, enabling controllable generation across multiple modalities without relying on registered data pairs or mask-annotated data from target modalities. Different colors represent distinct modalities.
@@ 1 Introduction
Medical image segmentation [43, 21, 33, 55] has achieved remarkable success, becoming a cornerstone of intelligent healthcare. However, challenges such as data privacy, modality complexity, and the high cost of segmentation annotations significantly restrict its development towards unannotated modalities. As depicted in Figure 1 (left), these limitations are especially acute in MRI, which exhibits significant discrepancies across modalities.
Magnetic Resonance Imaging (MRI) is a non-invasive imaging technique widely recognized for its high-quality and radiation-free characteristics. However, MRI scans are often expensive, time-consuming, and sensitive to motion. Additionally, there is significant variability across different MRI modalities, and even within the same modality, differences in machines and scanning parameters can lead to substantial discrepancies in images. The complexity and variability of MRI, coupled with high annotation costs, lead to a scarcity of registered data pairs and segmentation annotations, posing significant challenges to the development of robust segmentation models. In this paper, we aim to leverage the progress in generative models, specifically, diffusion models [18], to facilitate the training of segmentation models on unannotated MRI modalities.
Generally speaking, there are three challenges when applying the generative models to MRI data synthesis: (i) paradigm: as presented in Figure 1 (right), previous models [13, 3] primarily focus on generating data for modalities with annotations. This approach limits their scalability to under-annotated or unannotated MRI modalities; (ii) data: the scarcity of MR images and fine-grained annotations makes it challenging to develop general and controllable generative models; (iii) controllability: to better benefit downstream tasks with synthetic data, generative models must enable controllable generation, guided by conditions such as text prompts or segmentation masks, to produce training samples for downstream tasks.
@@
To tackle the above challenges, we collect and curate a large-scale radiology image-text dataset, MedGen-1M, featuring CT and MRI of various modalities sourced from the Internet and open-source data. Our dataset comprises 1.2 million 2D slices with modality labels, attributes, region, and organ information, with a subset also containing organ mask annotations. The extensive image-text pairs across diverse modalities provide a foundation for training a general MRI generation model, and the available mask annotations further enable more controllable generation.
@@
Building on this foundation, this paper introduces MRGen, a diffusion-based data engine for controllable MRI synthesis, particularly, for modalities with no segmentation masks ever available. At training time, we employ a two-stage strategy: (i) text-guided pretraining on diverse, large-scale radiology image-text pair data, enabling the model to generate images across various modalities based on text descriptions; and (ii) mask-conditioned finetuning on data with mask annotations, facilitating controllable generation based on organ masks. Consequently, such a strategy even allows MRGen to extend its controllable generation abilities towards modalities that do not have segmentation annotations available. During inference, MRGen takes text prompts and organ masks as inputs, producing MR images aligned with the given modality, region, and organs. With such synthetic samples, we can thus facilitate MRI segmentation tools towards modalities without manual annotations.
@@
Our contributions in this paper can be summarized as follows: (i) we establish a new paradigm for generative model applications in medical analysis, i.e., synthesizing data of unannotated modalities to serve as training samples for downstream segmentation models; (ii) we curate a large-scale radiology image-text dataset, MedGen-1M, which features detailed modality labels, attributes, regions, organ information, and a subset with organ mask annotations, providing a foundation for building general and controllable image generation models in the radiology domain; (iii) we develop MRGen, a diffusion-based data engine capable of generation conditioned on text prompts and masks, synthesizing MR images of various annotation-scarce modalities; (iv) we conduct extensive experiments across diverse modalities, demonstrating that MRGen can controllably synthesize high-quality MR images, boosting segmentation performance on unannotated modalities.
@@ 2 Related works
Generative models have been a research focus in computer vision for years, with GANs [12] and diffusion models [18, 45] leading the advancements. These models have found extensive applications across various tasks, including text-to-image generation [42, 24, 38, 48], image-to-image translation [4, 22, 58], artistic creation [30, 51, 44], and even challenging video generation [10, 19]. Notably, CycleGAN [58] employs cycle-consistency loss to facilitate image translation with unpaired data, while the Stable Diffusion series [42, 39, 11] efficiently produces high-resolution images in latent space, earning broad recognition.
Medical image synthesis aims to leverage generative models to improve medical image analysis models or tackle challenges such as data scarcity and privacy concerns [28]. Prior works train generative models on X-ray [3], CT [13], and brain MRI [8, 36], while methods like DiffTumor [5] and FreeTumor [49] focus specifically on generating tumor images to improve tumor segmentation.
Existing works [3, 13] primarily focus on in-domain augmentation within mask-annotated training modalities, and often struggle to generalize to unseen and unannotated modalities. To this end, this paper explores a novel paradigm for medical image generation, producing samples for modalities lacking mask annotations, to train MRI segmentation models towards these challenging scenarios.
@@
Medical image segmentation has been a long-standing research topic, with various architectures being proposed [43, 21, 15, 56, 34]. Recently, inspired by SAM [27, 41], large-scale interactive segmentation models [33, 55, 9] have been developed for medical image analysis. However, the heterogeneity of MRI challenges the generalization of current segmentation models, which struggle with varying intensity distributions among diverse modalities. Existing methods generally address this problem by learning domain-invariant content, which either requires mixing training data from multiple domains, or relies on deliberately designed data augmentations strategies [57, 20, 46, 37, 50, 6]. Despite recent efforts in annotating segmentation masks for extensive MRI data [7, 14], building datasets with wide-ranging modalities, scanners and protocols is time-consuming and labor-intensive. In this paper, we explore controllable generative models to synthesize mask-conditioned MR images for any segmentation model, without tedious hyperparameter selection in data augmentation.
3 Data Curation
To train our proposed data engine, we first collect and curate a large-scale radiology dataset, termed MedGen-1M, containing CT and MR images with rich modality information, clinical attributes, and mask annotations. In the following, we elaborate on our data processing pipeline and present the statistics of the curated dataset.
Data collection. We first collect a large amount of abdominal CT and MR images from the Radiopaedia111radiopaeida.org website, licensed under CC BY-NC-SA 3.0222To obtain the data, readers must first get permission from Radiopaedia, then we will share the download link.. This portion of data comprises a wide variety of modalities, serving as the large-scale image-text pairs ($\{(\mathcal{I},\mathcal{T})\}$), with each sample containing an image and its modality label. As listed in Table 1, we also integrate CT and MR data with modality labels and abdominal organ mask annotations from several open-source datasets, to construct data triplets ($\{(\mathcal{I},\mathcal{T},\mathcal{M})\}$), which will be further explained in Section 4.
Automatic annotations. Given the wide variability in abdominal imaging, such as differences between the upper abdominal region and the pelvic region, relying solely on modality labels may not sufficiently distinguish these images. Thus, we divide the abdomen into six regions, including: Upper Thoracic Region, Middle Thoracic Region, Lower Thoracic Region, Upper Abdominal Region, Lower Abdominal Region, and Pelvic Region. We then use off-the-shelf BiomedCLIP [52] to categorize all 2D slices into these six categories to serve as region information.
Moreover, considering that even medical-specific text encoders may not clearly distinguish fine-grained correlations and differences between modalities like T1 and T2 [47, 53], we employ GPT-4 [1] to map all modality labels to free-text modality attributes that describe signal intensities of fat, muscle, and water. For example, T1 modality can be represented as fat high signal, muscle intermediate signal, water low signal. This aims to help the model understand the imaging characteristics among distinct modalities.
Discussion. After the data processing mentioned above, we assemble our MedGen-1M dataset. As shown in Table 1, the dataset consists of approximately 16K 3D volumes, covering CT and over a hundred MR modalities, totaling nearly 1.2M 2D slices. Each sample is paired with its corresponding modality labels, modality attributes, region, and organ information, with 190K samples featuring paired mask annotations. Compared to the existing datasets, the scale, modality diversity, and fine-grained annotations of MedGen-1M provide more information for training generative models tailored for medical imaging. More statistics will be detailed in Section B.2 of the Appendix.
Dataset Modality #Volumes #Slices #Masks
Radiopaedia-MR MR 5,414 205,039
Radiopaedia-CT CT 8,734 812,756
LiQA [31] MR 30 2,185 1,446
PanSeg [54] MR 766 33,360 13,779
PROMISE12 [29] MR 50 1,377 778
MSD-Prostate [2] MR 64 1,204 366
CHAOS-MRI [25] MR 60 1,917 1,492
MSD-Liver [2] CT 131 58,638 19,163
AMOS22CT [23] CT 300 41,430 30,279
Total CT & MR 15,499 1,156,529 186,525
Table 1: Dataset Statistics of MedGen-1M.
@@ 3 Data Curation
Figure 2: Architecture Overview. Developing our MRGen involves three key steps: (a) Train an autoencoder on various images from dataset $\mathcal{D}_{u}$; (b) Train a text-guided generative model within the latent space, using image-text pairs across diverse modalities from $\mathcal{D}_{u}$, featuring modality, attributes, region, and organs information; (c) Train a mask condition controller jointly on the mask-annotated source-domain dataset $\mathcal{D}_{s}$ and unannotated target-domain dataset $\mathcal{D}_{t}$, enabling controllable generation based on both text prompts and masks.
@@ 4 Method
We first formulate the problem in Section 4.1, followed by a detailed description of the proposed MRGen architecture in Section 4.2; then we elaborate on the training details of our model in Section 4.3; lastly, we present the procedure of synthesizing and filtering samples with our data engine for downstream segmentation tasks in Section 4.4.
@@ 4.1 Problem Formulation
In this paper, we focus on developing a diffusion-based controllable data engine for medical imaging, particularly for synthesizing data across various modalities with no mask annotations available. Specifically, our proposed MRGen ($\Phi_{\mathrm{MRGen}}$) generates the radiology image ($\mathcal{I}$), conditioned on a text prompt ($\mathcal{T}$) and the organs mask ($\mathcal{M}$), i.e., $\mathcal{I}=\Phi_{\mathrm{MRGen}}(\mathcal{T},\mathcal{M};\Theta,\Theta_{c})$, where $\Theta$ and $\Theta_{c}$ refer to the trainable parameters of the foundation generative model and the mask condition controller, respectively.
As described in Section 3, our large-scale dataset, i.e., $\mathcal{D}=\{\mathcal{D}_{u},\mathcal{D}_{s},\mathcal{D}_{t}\}$, encompasses three types of data, supporting the development of the proposed data engine. Concretely, the training process is carried out in two phases: (i) we pretrain a text-guided generative model on the image-text data ($\mathcal{D}_{u}=\{(\mathcal{I}_{u},\mathcal{T}_{u})\}$), covering diverse modalities; and (ii) we then train a condition controller jointly on image-text pairs with mask annotations from source-domain dataset ($\mathcal{D}_{s}=\{(\mathcal{I}_{s},\mathcal{T}_{s},\mathcal{M}_{s})\}$), and image-text pairs from the target-domain dataset ($\mathcal{D}_{t}=\{(\mathcal{I}_{t},\mathcal{T}_{t},\varnothing)\}$). The ultimate goal is to generate target-domain images ($\mathcal{I^{\prime}}_{t}$), conditioned on text prompts ($\mathcal{T^{\prime}}_{t}$) and masks ($\mathcal{M^{\prime}}_{t}$), to serve as training samples for segmentation models.
Relations to existing tasks. This paper makes two key advancements over existing methods: (i) conventional generative augmentation methods are confined to generating samples within mask-annotated training modalities [13, 3, 8]. Conversely, our model can synthesize data for modalities lacking mask annotations, conditioned on text prompts and masks, improving segmentation models on these challenging scenarios; (ii) unlike image translation works [58, 22, 36] that typically require registered data pairs for training or are limited to specific conversions, our model provides flexible and controllable generation with no need for registered data or mask-annotated data of the target modality.

### v2

# MRGen: Segmentation Data Engine For Underrepresented MRI Modalities

@@ Abstract
Training medical image segmentation models for rare yet clinically significant imaging modalities is challenging due to the scarcity of annotated data, and manual mask annotations can be costly and labor-intensive to acquire. This paper investigates leveraging generative models to synthesize training data, to train segmentation models for underrepresented modalities, particularly on annotation-scarce MRI. Concretely, our contributions are threefold: (i) we introduce MRGen-DB, a large-scale radiology image-text dataset comprising extensive samples with rich metadata, including modality labels, attributes, regions, and organs information, with a subset having pixelwise mask annotations; (ii) we present MRGen, a diffusion-based data engine for controllable medical image synthesis, conditioned on text prompts and segmentation masks. MRGen can generate realistic images for diverse MRI modalities lacking mask annotations, facilitating segmentation training in low-source domains; (iii) extensive experiments across multiple modalities demonstrate that MRGen significantly improves segmentation performance on unannotated modalities by providing high-quality synthetic data. We believe that our method bridges a critical gap in medical image analysis, extending segmentation capabilities to scenarios that are challenging to acquire manual annotations. The codes, models, and data will be publicly available at https://haoningwu3639.github.io/MRGen/.
Figure 1: Motivations and Overview. Left: The heterogeneity of MRI modalities challenges the generalization of segmentation models. Our proposed data engine, MRGen, overcomes this by controllably synthesizing training data for segmentation models. Right: (a) Prior generative models are restricted to data augmentation for well-annotated modalities; (b) Image translation typically requires registered data pairs (dashed lines), and is limited to specific modality conversions; (c) MRGen enables controllable generation across diverse modalities, creating data for training segmentation models towards underrepresented modalities. Distinct colors represent different modalities.
@@ 1 Introduction
Medical image segmentation [50, 23, 39, 5] has shown remarkable success by training on extensive manual annotations, becoming a cornerstone of intelligent healthcare systems. However, developing models for underrepresented imaging modalities remains challenging, due to data privacy, modality complexity, and high cost of manual mask annotations [8, 16], especially for rare yet clinically important modalities, for example, Magnetic Resonance Imaging (MRI). Despite being non-invasive and radiation-free, MRI scanning is expensive and exhibits substantial variability across modalities and scanning protocols [43]. This lack of standardization and numerous hyperparameters fragment the already limited dataset, challenging the development of robust segmentation models, as illustrated in Figure 1 (left).
In this paper, we investigate the potential of generative models, particularly diffusion models, to synthesize MRI data for training segmentation models on underrepresented modalities. While generative models offer a promising solution, they face unique challenges: (i) data availability remains a significant obstacle. Existing approaches have primarily focused on data augmentation for well-annotated modalities such as X-ray [3] and CT [14, 15], as depicted in Figure 1 (right). However, MRI data is relatively scarce, and highly diverse across modalities, making it less explored for generative modeling; (ii) controllability is critical to facilitate downstream tasks such as segmentation. Thus, generative models must enable controllable synthesis based on conditions such as texts, masks, or both. Yet, prior works [55, 9, 42, 60, 32] cannot simultaneously support both conditions, limiting their abilitiy to control the generated modalities, regions, and organs effectively.
@@
Our first contribution is the collection of a large-scale, high-quality radiology image-text dataset, MRGen-DB (short for “Database for MRI Generation”), which includes MRI scans across various modalities sourced from the Internet and open-source repositories. The dataset consists of nearly 250,000 2D slices, enriched with detailed annotations such as modality labels, attributes, region, and organ information, with a subset providing organ masks. This extensive collection of image-text pairs across diverse modalities forms a robust foundation for training a general MRI generation model, while the availability of mask annotations facilitates more controllable and targeted synthesis.
@@
For controllable data generation, we present MRGen, a diffusion-based data engine for MRI synthesis, that supports conditioning on both text prompts and segmentation masks. We employ a two-stage training strategy: (i) text-guided pretraining on diverse, large-scale image-text pairs, enabling the model to synthesize images across various modalities guided by templated text descriptions; and (ii) mask-conditioned finetuning on mask-annotated data, facilitating controllable generation based on organ masks. Consequently, such a two-stage strategy allows MRGen to extend its controllable generation abilities towards modalities that originally do not have segmentation annotations available, thereby enables training segmentation models for these underrepresented modalities with synthetic data.
@@
Overall, our contributions can be summarized as follows: (i) we explore the use of generative models for MRI synthesis across annotation-scarce modalities, facilitating training segmentation models for underrepresented modalities; (ii) we curate MRGen-DB, a large-scale radiology image-text dataset, which features detailed modality labels, attributes, regions, and organs information, with a subset of organ mask annotations, providing a robust foundation for medical generative modeling; (iii) we develop MRGen, a diffusion-based data engine capable of controllable generation, conditioned on templated text prompts and segmentation masks; (iv) we conduct extensive experiments across diverse modalities, demonstrating that MRGen can controllably generate high-quality MR images, improving ‘zero-shot’ segmentation performance for unannotated modalities. To our knowledge, this work introduces the first open-source dataset curated for medical image generation, and the first general medical generative model tailored for annotation-scarce MRI modalities, offering a novel solution to address the scarcity of medical data and annotations.
@@ 2 Related works
Generative models have been a research focus in computer vision for years, with GANs [13] and diffusion models [20, 53] leading the advancements. These models have found extensive applications across various tasks, including text-to-image generation [49, 25, 45, 57], image-to-image translation [4, 24, 69], artistic creation [35, 61, 51], and even challenging video generation [11, 21]. Notably, CycleGAN [69] employs cycle-consistency loss to facilitate image translation with unpaired data, while Stable Diffusion series [49, 47, 12] efficiently produces high-resolution images in latent space, earning broad recognition.
Medical image synthesis aims to leverage generative models to tackle challenges such as data scarcity [37, 32], biases [33], and privacy concerns [31]. Prior works primarily focus on X-ray [3], CT [14, 15], and brain MRI [9, 42, 64], with approaches like DiffTumor [6] and FreeTumor [58] specifically targeting tumor generation to boost tumor segmentation. While these methods have proven effectiveness in data augmentation within well-annotated training modalities and regions, they still struggle to generalize to modalities lacking manual mask annotations. To this end, this paper investigates adopting generative models to facilitate more robust segmentation models with high-quality synthetic data of annotation-scarce modalities.
@@
Medical image segmentation has been a long-standing research topic, with various architectures proposed [50, 23, 17, 67, 40]. Recently, inspired by SAM [30, 48], large-scale segmentation models [39, 66, 10] have been developed. However, the heterogeneity of MRI challenges the generalization of existing models, which struggle with intensity variations among diverse modalities. Existing methods attempt to address this with image translation [52, 46, 28] or relying on deliberately designed augmentation strategies to learn domain-invariant content [68, 22, 54, 44, 59, 7]. In this paper, we explore controllable generative models to synthesize data for segmentation training, particularly towards underrepresented modalities lacking mask annotations, thus resembling a ‘zero-shot’ segmentation scenario.
@@
Figure 2: Architecture Overview. Developing our MRGen involves three key steps: (a) Train an autoencoder on various images from dataset $\mathcal{D}_{u}$; (b) Train a text-guided generative model within the latent space, using image-text pairs across diverse modalities from MRGen-DB, featuring modality, attributes, region, and organs information; (c) Train a mask condition controller jointly on image-text pairs with and without mask annotations, enabling controllable generation based on both text prompts and masks.
@@ 3 Method
Here, we start to formulate the problem of interest in Sec. 3.1, followed by a detailed description of the dataset curation procedure in Sec. 3.2. Later, we elaborate on the proposed MRGen architecture and the training details of our model in Sec. 3.3 and Sec. 3.4, respectively. Lastly, we present the procedure of synthesizing and filtering samples for downstream segmentation tasks in Sec. 3.5.
@@ 3.1 Problem Formulation
Given a text prompt ($\mathcal{T}$) describing modality, region, and organs, along with the organs mask ($\mathcal{M}$), our proposed MRGen ($\Phi_{\mathrm{MRGen}}$) enables to generate the MR image ($\mathcal{I}$):
$\mathcal{I}=\Phi_{\mathrm{MRGen}}(\mathcal{T},\mathcal{M};\Theta,\Theta_{c})$
where $\Theta$ and $\Theta_{c}$ refer to parameters of the generative model and the mask condition controller, respectively. Developing such a controllable data engine, thus enables synthesizing high-quality data to train segmentation models for the challenging ‘zero-shot’ scenario.
Relations to existing tasks. Numerous studies have proven the effectiveness of generative models for data augmentation [55] on well-annotated modalities and regions, such as CT [15, 14], X-ray [3], and brain MRI [9, 42], however, this is not the focus of our work. Instead, we target the more challenging scenario of synthesizing MR images for scarce and underrepresented modalities where no manual mask annotations are available. While image translation methods [69, 27, 52, 46] offer a potential alternative by translating richly annotated data to underrepresented modalities, these approaches often require registered data for training or are limited to specific modality conversions. In contrast, our proposed MRGen framework offers a flexible and controllable generation pipeline, enabling the synthesis of complex abdominal MRI data across diverse modalities, even in the absence of mask annotations.
Dataset # Volumes # Slices # Masks
Radiopaedia-MRI 5,414 205,039
PanSeg [65] 766 33,360 13,779
MSD-Prostate [2] 64 1,204 366
CHAOS-MRI [26] 60 1,917 1,492
PROMISE12 [34] 50 1,377 778
LiQA [36] 30 2,185 1,446
Total 6,384 245,082 17,861
Table 1: Dataset Statistics of MRGen-DB.
3.2 Dataset Curation
The scarcity of MR images with comprehensive text descriptions and mask annotations poses challenges for training generative models. To tackle this limitation, we present MRGen-DB (short for “Database for MRI Generation”), a meticulously curated large-scale radiology dataset featuring diverse MR images enriched with modality information, detailed clinical attributes, and precise mask annotations. Below, we detail our data processing pipeline and provide comprehensive dataset statistics.
Data collection. The primary source of our dataset is abdominal MR images obtained from Radiopaedia [^1] Radiopaedia.org., licensed under CC BY-NC-SA 3.0222Access to this dataset requires prior permission from Radiopaedia, after which we will provide the download link.. This portion of data includes a diverse array of imaging modalities, forming extensive image-text pairs suitable for training text-guided generative models. Each sample consists of an MR image and its corresponding free-text modality label.
To enhance the dataset’s coverage and utility, we also augment it with abdominal MRI data from multiple open-source repositories. These additional data sources include modality labels and organ-specific mask annotations, forming comprehensive data triplets ($\{\mathcal{I},\mathcal{T},\mathcal{M}\}$). This augmentation enables more sophisticated controllable generation guided by both textual descriptions and anatomical masks, broadening the dataset’s applicability.
Automatic annotations. Abdominal imaging is highly variable, exhibiting significant differences across anatomical regions, such as the Upper Abdominal Region and the Pelvic Region. Relying solely on modality labels is insufficient to differentiate these distinctions. To address this, we divide the abdomen into six anatomical regions: Upper Thoracic Region, Middle Thoracic Region, Lower Thoracic Region, Upper Abdominal Region, Lower Abdominal Region, and Pelvic Region. Using the pre-trained BiomedCLIP model [62], we automatically categorize all 2D slices into these regions. To maintain annotation quality, slices with low confidence scores ($<40\%$) are intentionally left unlabeled. This process enriches the dataset with detailed region-specific information.
Distinguishing fine-grained modality differences, such as between T1 and T2, presents additional challenges, even for advanced medical-specific text encoders [56, 63]. To overcome this, we employ GPT-4 [1] to map modality labels into free-text attributes that describe the signal intensities of specific tissues, including fat, muscle, and water. For example, the T1 modality can be represented as: fat high signal, muscle intermediate signal, water low signal. These detailed descriptions enable the model to understand and differentiate imaging characteristics across modalities.
To ensure the reliability of the automatic annotations, we have uniformly sampled and manually verified a subset of the data. Specifically, 2% of the region annotations and 20% of the modality attributes have been reviewed, achieving high accuracies of 95.33% and 91.67%, respectively. This verification step ensures the high quality of the dataset and strengthens its applicability for downstream tasks.
Discussion. After completing the data processing above, we assemble the MRGen-DB dataset, which includes approximately 6,000 3D volumes spanning over 100 distinct free-text MR modalities, totaling nearly 250,000 2D slices, as presented in Table 1. Each sample is paired with its modality label, attributes, region, and organ information, with about 18,000 samples featuring mask annotations. The scale, diversity, and fine-grained annotations of MRGen-DB provide sufficient information for training generative models tailored for MR images. More statistics are provided in Sec. B.2.